A silver coin was purchased for $430 in 1985.The market value for silver declined to the point that in 2007 the coin was worth $200. What was the average annual rate of depreciation of the silver coin?
Answer by stanbon(75887)

A silver coin was purchased for $430 in 1985.The market value for silver declined to the point that in 2007 the coin was worth $200. What was the average annual rate of depreciation of the silver coin?
-------------------
(200-430)/(2007-1985) = -230/22 = -$10.4545 per year
======================
Cheers,
Stan H.